OpenStudy (anonymous):

Solve by completing the square: x^2-12x-5=0

OpenStudy (anonymous):

I know that you add 5 so it would be: x^2-12x+_c_=5 Then I dont know how to find C?

OpenStudy (zale101):

x^2-12x-5=0 x^2-12x=5 Move the number(without the variable) over to the other side.

OpenStudy (zale101):

take half of 12x, and square it

OpenStudy (anonymous):

Then to find C I half 12 then square it and that would be 36 right?

OpenStudy (anonymous):

x^2-12x+36=41

OpenStudy (anonymous):

SO: (x-6)^2=square root of + or-41

OpenStudy (jdoe0001):

hmmm usually you'd want to group the "x" terms or like-terms

OpenStudy (zale101):

36 + -12x + x^2 = 41 Factor a perfect square on the left side (x + -6)(x + -6) = 41

OpenStudy (zale101):

i think you're on the hang of it